If given 285mL of a diluted glucose solution 0.75% (w/v) that was made by adding 35mL of a stock glucose solution and 250mL of water, what was the original concentration of the stock glucose solution expressed as grams/100mL or % (w/v)
Answer #1

According to the law of equilibrium C1V1= C2V2

C1= concentration of the stock solution. We need to find out

V1 - the volume of stock solution that is used to make the working solution. 35 ml

C2 - concentration of the working solution= 0.75 % (w/v)

V2 - Final volume = 285 ml

35 * C1=0.75 * 285

C1 = 6.10 (w/v%)

or 6.10 gm of glucose per 100ml water

So the concentration of the stock solution was 6.10% (w/v)